# TexVerse: A Universe of 3D Objects with High-Resolution Textures

### v1

@@ 1 Introduction
To fill this gap, we introduce TexVerse, a large-scale 3D dataset featuring high-resolution textures. Sketchfab hosts approximately 1.6 million freely downloadable 3D models. We curate TexVerse by filtering models with texture resolutions of at least 1024 pixels, excluding those tagged or described with terms related to “NoAI” and retaining only models licensed under distributable Creative Commons licenses. TexVerse encompasses 858,669 distinct high-resolution 3D models, of which 158,518 incorporate PBR materials, all standardized in the .glb format. Additionally, for each model, we also collect all of its high-resolution variants (e.g., the 4096 and 1024 versions of a model with a maximum resolution of 8192), yielding a total of 1,659,097 3D instances. For the rigged and animated categories of models, we further obtain the original user-uploaded file format to prevent the loss of skeletons and animations during the format conversion of Sketchfab. These are named as the TexVerse-Skeleton and TexVerse-Animation datasets, comprising 69,138 rigged models and 54,430 animated models, respectively, all with high-resolution textures. Furthermore, we provide 856,312 detailed annotations to the model generated by GPT-5 [2], which encompass the general description, structural composition and detailed characteristics. We believe TexVerse will drive the community forward in areas such as high-resolution texture generation, PBR material synthesis, animation, and a wide range of 3D vision and graphics tasks.
@@ 3 TexVerse
TexVerse is a large-scale 3D dataset featuring high-resolution textures. The objects are sourced from Sketchfab, an online 3D marketplace where users can upload and share models for both free and commercial use. We conducted a comprehensive survey of Sketchfab, identifying approximately 1.6 million freely downloadable 3D models uploaded between 2012 and 2025. Using metadata provided by Sketchfab, we first filtered for high-resolution textured models with texture resolutions of at least 1024 pixels, excluding models tagged or described with terms related to “NoAI”. Then, we obtained models which under the distributable Creative Commons license using Sketchfab’s API. The resulting dataset, formatted in the .glb format, comprises 858,669 unique high-resolution textured 3D objects across various resolution levels. For each model, we also collected all of its high-resolution variants (e.g., the 4096 and 1024 versions of a model with a maximum resolution of 8192), yielding a total of 1,659,097 3D instances.
@@ 3.1 Metadata
For each object in the TexVerse, we provide metadata extracted from the information supplied by its creator upon uploading to Sketchfab: uid, user-name, tags, categories, thumbnail-url, vertex-count, face-count, max-texture, pbr-type, is-rigged, animation-count, license.

### v2

@@ 1 Introduction
To fill this gap, we introduce TexVerse, a large-scale 3D dataset featuring high-resolution textures. Sketchfab hosts approximately 1.6 million freely downloadable 3D models. We curate TexVerse by filtering models with texture resolutions of at least 1024 pixels, excluding those tagged or described with terms related to “NoAI” and retaining only models licensed under distributable Creative Commons licenses. TexVerse encompasses 858,669 distinct high-resolution 3D models, of which 158,518 incorporate PBR materials, all standardized in the .glb format. Additionally, for each model, we also collect all of its high-resolution variants (e.g., the 4096 and 1024 versions of a model with a maximum resolution of 8192), yielding a total of 1,661,101 3D instances. For the rigged and animated categories of models, we further obtain the original user-uploaded file format to prevent the loss of skeletons and animations during the format conversion of Sketchfab. These are named as the TexVerse-Skeleton and TexVerse-Animation datasets, comprising 69,138 rigged models and 54,430 animated models, respectively, all with high-resolution textures. Furthermore, we provide 856,312 detailed annotations to the model generated by GPT-5 [2], which encompass the general description, structural composition and detailed characteristics. We believe TexVerse will drive the community forward in areas such as high-resolution texture generation, PBR material synthesis, animation, and a wide range of 3D vision and graphics tasks.
@@ 3 TexVerse
TexVerse is a large-scale 3D dataset featuring high-resolution textures. The objects are sourced from Sketchfab, an online 3D marketplace where users can upload and share models for both free and commercial use. We conducted a comprehensive survey of Sketchfab, identifying approximately 1.6 million freely downloadable 3D models uploaded between 2012 and 2025. Using metadata provided by Sketchfab, we first filtered for high-resolution textured models with texture resolutions of at least 1024 pixels, excluding models tagged or described with terms related to “NoAI”. Then, we obtained models which under the distributable Creative Commons license using Sketchfab’s API. The resulting dataset, formatted in the .glb format, comprises 858,669 unique high-resolution textured 3D objects across various resolution levels. For each model, we also collected all of its high-resolution variants (e.g., the 4096 and 1024 versions of a model with a maximum resolution of 8192), yielding a total of 1,661,101 3D instances.
@@ 3.1 Metadata
For each object in the TexVerse, we provide metadata extracted from the information supplied by its creator upon uploading to Sketchfab: uid, name, description, user-name, tags, categories, thumbnail-url, vertex-count, face-count, max-texture, pbr-type, is-rigged, animation-count, license.